### v2

# On the maximum $F$-free induced subgraphs in $K_{t}$-free graphs

Abstract
For graphs $F$ and $H$, let $f_{F,H}(n)$ be the minimum possible size of a maximum $F$-free induced subgraph in an $n$-vertex $H$-free graph. This notion generalizes the Ramsey function and the Erdős–Rogers function. Establishing a container lemma for the $F$-free subgraphs, we give a general upper bound on $f_{F,H}(n)$, assuming the existence of certain locally dense $H$-free graphs. In particular, we prove that for every graph $F$ with $\mathrm{ex}(m,F)=O(m^{1+\alpha})$, where $\alpha\in[0,1/2)$, we have
$f_{F,K_{3}}(n)=O\left(n^{\frac{1}{2-\alpha}}\left(\log n\right)^{\frac{3}{2-\alpha}}\right)\quad\textrm{and}\quad f_{F,K_{4}}(n)=O\left(n^{\frac{1}{3-2\alpha}}\left(\log n\right)^{\frac{6}{3-2\alpha}}\right).$
For the cases where $F$ is a complete multipartite graph, letting $s=\sum_{i=1}^{r}s_{i}$, we prove that
$f_{K_{s_{1},\ldots,s_{r}},K_{r+2}}(n)=O\left(n^{\frac{2s-3}{4s-5}}(\log n)^{3}\right).$
We also make an observation which improves the bounds of $\mathrm{ex}(G(n,p),C_{4})$ by a polylogarithmic factor.
1 Introduction
For graphs $F$ and $G$, we say that a subset $S\subseteq V(G)$ is $F$-independent if the induced subgraph $G[S]$ does not contain $F$ as a subgraph. The $F$-independence number $\alpha_{F}(G)$ of $G$ is the maximum size of an $F$-independent subset in $G$. For graphs $F$ and $H$, let $f_{F,H}(n)$ be the minimum possible value of $\alpha_{F}(G)$ over all $n$-vertex $H$-free graphs $G$. If $F=K_{i}$ and $H=K_{j}$, i.e., they are complete graphs, then we simply write $f_{i,j}(n)$ for $f_{K_{i},K_{j}}(n)$.
In 1962, Erdős and Rogers [10] initiated the study of $f_{s,t}(n)$, which is now commonly referred to as the Erdős–Rogers function. Note that the case $s=2$ is related to Ramsey theory [23]: the Ramsey number $r(s,t)$ is at least $n$ if and only if $f_{2,t}(n)<s$. For example, by the bound $r(3,t)=\Theta(t^{2}/\log t)$ [1, 13] and the recent breakthrough $\Omega(t^{3}/(\log t)^{4})=r(4,t)=O(t^{3}/(\log t)^{2})$ [1, 18], we have
$f_{2,3}(n)=\Theta\left(n^{\frac{1}{2}}(\log n)^{\frac{1}{2}}\right)\quad\textrm{and}\quad\Omega\left(n^{\frac{1}{3}}(\log n)^{\frac{2}{3}}\right)=f_{2,4}(n)=O\left(n^{\frac{1}{3}}(\log n)^{\frac{4}{3}}\right).$ (1)
For $s>2$, Erdős and Rogers [10] proved that $f_{s,s+1}(n)\leqslant n^{1-\Omega_{s}(1)}$. In the following decades, there were many papers on this topic and significant improvements have been achieved for various pairs of $(s,t)$. The current best bound [8, 20] on $f_{s,s+1}(n)$ is
$\Omega\left((n\log n/\log\log n)^{\frac{1}{2}}\right)=f_{s,s+1}(n)=O\left(n^{\frac{1}{2}}\log n\right),$
and for $f_{s,s+2}(n)$, the best known bound [25, 12] is
$n^{\frac{1}{2}-\frac{1}{6s-6}}(\log n)^{\Omega(1)}=f_{s,s+2}(n)=O\left(n^{\frac{2s-3}{4s-5}}(\log n)^{3}\right).$
There are also many results on general $(s,t)$, see [6, 15, 16, 26, 25]. For more results about the Erdős–Rogers function, we refer the reader to the recent paper [12] and the citations within.
In this paper, we investigate some general cases of $f_{F,H}(n)$, where $F$ is not necessarily a complete graph. This problem was suggested by Dhruv Mubayi, when he visited UIUC in Fall 2023. By definition, we have $f_{F_{1},H}(n)\leqslant f_{F_{2},H}(n)$ whenever $F_{1}\subseteq F_{2}$, so $f_{2,H}(n)$ serves as a lower bound of $f_{F,H}(n)$ for every $F$ containing at least one edge. For the upper bound [^1] We are grateful to Jacques Verstraëte for pointing out this to us., note that every $F$-free subgraph in any graph $G$ has size less than $r(F,\alpha(G)+1)$. Hence, every upper bound on $r(F,t)$ and $f_{2,H}(n)$ offers an upper bound for $f_{F,H}(n)$. For example, using (1) and $r(C_{\ell},t)=O(t^{1+1/(\lceil\ell/2\rceil-1)})$ [9], we deduce that for every $k\geqslant 2$,
$\Omega\left(n^{\frac{1}{2}}(\log n)^{\frac{1}{2}}\right)=f_{C_{2k},K_{3}}(n)=O\left(n^{\frac{k}{2k-2}}\left(\log n\right)^{\frac{k}{2k-2}}\right),$
and
$\Omega\left(n^{\frac{1}{3}}(\log n)^{\frac{2}{3}}\right)=f_{C_{2k},K_{4}}(n)=O\left(n^{\frac{k}{3k-3}}\left(\log n\right)^{\frac{4k}{3k-3}}\right),$
and the same bounds hold for $f_{C_{2k-1},K_{3}}(n)$ and $f_{C_{2k-1},K_{4}}(n)$.
Our first result is an improvement on the upper bounds of $f_{F,K_{3}}(n)$ and $f_{F,K_{4}}(n)$ where $F$ has a small Turán number. Recall that the Turán number $\mathrm{ex}(n,F)$ of $F$ is the maximum number of edges in an $n$-vertex $F$-free graph.
Theorem 1.
For every graph $F$ with $\mathrm{ex}(b,F)\leqslant O(b^{1+\alpha})$, where $\alpha\in[0,1/2)$, we have
$f_{F,K_{3}}(n)=O\left(n^{\frac{1}{2-\alpha}}\left(\log n\right)^{\frac{3}{2-\alpha}}\right)\quad\quad\textrm{and}\quad\quad f_{F,K_{4}}(n)=O\left(n^{\frac{1}{3-2\alpha}}\left(\log n\right)^{\frac{6}{3-2\alpha}}\right).$
In particular, by the Bondy–Simonovits theorem [7], we have $\mathrm{ex}(b,C_{2k})=O(b^{1+1/k})$ for every integer $k\geqslant 2$. Therefore, we have the following corollary.
Corollary 2.
For every integer $k\geqslant 3$, we have
$f_{C_{2k},K_{3}}(n)=O\left(n^{\frac{k}{2k-1}}\left(\log n\right)^{\frac{3k}{2k-1}}\right)\quad\quad\textrm{and}\quad\quad f_{C_{2k},K_{4}}(n)=O\left(n^{\frac{k}{3k-2}}\left(\log n\right)^{\frac{6k}{3k-2}}\right).$
The main tool for the proof of Theorem 1 is a container lemma (Lemma 8), which offers an upper bound on the number of $F$-independent sets of a certain size in a “locally dense” graph. Using this Container Lemma, we can actually prove the following much more general result.
Theorem 3.
For real numbers $\alpha$, $\beta$, $\gamma$, $\delta$, and $\theta$, where $10\leqslant\alpha<$ 1, $0<\beta,\delta,\gamma$, $10<\theta<$ 1, and $\beta<\theta(1-\alpha)$, and a graph $F$ with $\mathrm{ex}(b,F)\leqslant O(b^{1+\alpha})$, there exist positive real numbers $c$, $C$, and $N$ such that the following holds for every integer $n>N$. Let
$s\coloneqq\left\lceil n^{\frac{\beta}{1-\alpha}}(\log n)^{\frac{3}{1-\alpha}}\right\rceil.$
If $G$ is an $n$-vertex graph with $e(G[A])\geqslant|A|^{2}\cdot\delta n^{-\beta}$ for every $A\subseteq V(G)$ with $|A|>\gamma n^{\theta}$, then there exists $G^{\prime}\subseteq G$ with at least $m\coloneqq c{s}n^{1-\theta}$ vertices and
$\alpha_{F}(G^{\prime})\leqslant s\leqslant C\cdot m^{\frac{\beta}{(1-\alpha)(1-\theta)+\beta}}(\log m)^{\frac{3(1-\theta)}{(1-\alpha)(1-\theta)+\beta}}.$
We will derive Theorem 1 from Theorem 3 by letting $G$ be the pseudorandom $K_{3}$-free graph given by Alon [2] and the construction of certain $K_{4}$-free graphs given by Mattheus and Verstraëte [18], respectively.
Theorem 3 also implies many other results, such as the following theorem, which gives an upper bound on $f_{C_{2k},K_{t}}(n)$, assuming the existence of “perfectly” pseudorandom $K_{t}$-free graphs. Recall that an $(n,d,\lambda)$-graph is an $n$-vertex $d$-regular graph such that the absolute value of every eigenvalue of its adjacency matrix, but the largest one, is at most $\lambda$. For more background about $(n,d,\lambda)$-graphs, we refer the reader to the survey [17].
Theorem 4.
For every graph $F$ and integer $t\geqslant 3$, where $\mathrm{ex}(b,F)=O(b^{1+\alpha})$ for some $\alpha\in[0,(t-2)/(t-1))$, if for every sufficiently large $m$, there exists a $K_{t}$-free $(m,d,\lambda)$-graph $G_{m}$ with $d=\Theta(m^{1-1/(2t-3)})$ and $\lambda=O(d^{1/2})$, then we have
$f_{F,K_{t}}(n)=O\left(n^{\frac{1}{(1-\alpha)(t-2)+1}}(\log n)^{\frac{3(t-2)}{(1-\alpha)(t-2)+1}}\right).$
An interesting observation is that, while different $K_{4}$-free graphs are used, Theorems 1 and 4 give the same upper bound on $f_{F,K_{4}}(n)$ for every graph $F$ with $\mathrm{ex}(b,F)=O(b^{1+\alpha})$ for some $\alpha\in[0,1/2)$.
For $F=C_{4}$ and $H=K_{4}$, Theorems 3 and 4 fail to provide a reasonable bound on $f_{F,H}(n)$. Instead, we have the following slightly more general theorem. Denote by $K_{s_{1},\ldots,s_{r}}$ the complete $r$-partite graph with parts of sizes $s_{1},\ldots,s_{r}$, respectively.
Theorem 5.
For every integer $r\geqslant 2$ and positive integers $s_{1},\ldots,s_{r}$, letting $s\coloneqq\sum_{i=1}^{r}s_{i}$, we have
$f_{K_{s_{1},\ldots,s_{r}},K_{r+2}}(n)=O\left(n^{\frac{2s-3}{4s-5}}(\log n)^{3}\right).$
In particular, Theorem 5 implies
$f_{C_{4},K_{4}}(n)=O\left(n^{\frac{5}{11}}(\log n)^{3}\right)\quad\text{and ingeneral }\quad f_{K_{a,a},K_{4}}(n)=O\left(n^{\frac{1}{2}-\frac{1}{16a-10}}(\log n)^{3}\right).$
Our proof for Theorem 5 follows closely the recent proof of $f_{{r},{r+2}}(n)=O(n^{(2r-3)/(4r-5)}(\log n)^{3})$ by Janzer and Sudakov [12], which followed closely the work of Mattheus and Verstraëte [18]. We generalize their ideas to bound the number of copies of $K_{s_{1},\ldots,s_{r}}$ in the graphs based on the construction of $K_{4}$-free graphs given in [18].
Remark 6.
While working on this project, we learned that Mubayi and Verstraëte [21] proved that $f_{F,K_{3}}(n)=n^{1/2+o(1)}$ for every triangle-free graph $F$, and for every $K_{4}$-free graph $F$ containing a cycle, there is a constant $c_{F}>0$ such that $f_{F,K_{4}}(n)>n^{1/3+c_{F}+o(1)}$. In particular, their results imply (and somewhat show the tightness of some of our results) that $f_{C_{2k},K_{4}}(n)=n^{1/3+\Theta(1/k)}$.
We also have the following observation on the random Turán number of $C_{4}$. Recall that $\mathrm{ex}(G(n,p),C_{4})$ is the maximum number of edges in a $C_{4}$-free subgraph of the Erdős–Rényi graph $G(n,p)$.
Proposition 7.
If $p\geqslant n^{-1/3}(\log n)^{8/3}$, then with high probability, we have
$\mathrm{ex}(G(n,p),C_{4})=O\left(p^{1/2}n^{3/2}\right).$
Proposition 7 improves a result of Morris and Saxton [19], which gives the same upper bound under the assumption that $p\geqslant n^{-1/3}(\log n)^{4}$. See Section 4 for more discussion on this problem.
The rest of this paper is organized as follows. In Section 2, we prove our Container Lemma and use it to prove Theorems 31, and 4. In Section 3, we prove Theorem 5, and in Section 4, we prove Proposition 7.
Notation. For a graph $G$, a subset $S\subseteq V(G)$, and a vertex $v\in V(G)$, we write $N_{S}(v)$ for the set of neighbors of $v$ in $S$ and let $d_{S}(v)\coloneqq|N_{S}(v)|$ be the degree of $v$ in $S$. When $S=V(G)$, we write $N_{G}(v)$ for $N_{V(G)}(v)$ and $d_{G}(v)$ for $d_{V(G)}(v)$. All logarithms are in base $e$, except in Section 3, where they are in base $2$.
2 The Container Method
Our main tool for proving Theorem 3 is the following Container Lemma, whose proof is an extended variant of the graph container method. See [4] for a prior similar application of the method.
Lemma 8 (Container Lemma).
For every graph $F$ and positive real numbers $d,r$, if $G$ is an $n$-vertex graph, where $e(G[A])\geqslant d\cdot|A|^{2}$ for every $A\subseteq V(G)$ with $|A|>r$, then for every $F$-independent subset $S\subseteq V(G)$ with $\mathrm{ex}(|S|,F)\cdot d\geqslant\log n$, there exist $S_{h}$, $T$, $C\subseteq V(G)$, where $C$ depends only on $S_{h}$, $T$, and $|S|$, such that
  1. (i)
    $(S_{h}\cup T)\subseteq S$ and $S\setminus(S_{h}\cup T)\subseteq C$,
  2. (ii)
    $|S_{h}|\leqslant 2\cdot\left(\frac{\mathrm{ex}(|S|,\,F)\log n}{d}\right)^{1/2}$,
  3. (iii)
    $|T|\leqslant 32\cdot\left(\frac{\mathrm{ex}(|S|,\,F)\log n}{d}\right)^{1/2}$,
  4. (iv)
    $|C|\leqslant r$.
Proof.
Choose $S\subseteq V(G)$ such that $G[S]$ is $F$-free and $\mathrm{ex}(|S|,F)\cdot d\geqslant\log n$. Letting $s\coloneqq|S|$, we have $e(G[S])\leqslant\mathrm{ex}(s,F)$. We use the following algorithm which takes $S$ and outputs the sets $S_{h}$, $T$, and $C$.
Let $\Delta\coloneqq(\mathrm{ex}(s,F)\cdot d/\log n)^{1/2}\geqslant 1$. Let $S_{h}\coloneqq\{v\in S:d_{S}(v)\geqslant\Delta\},$ and $S_{\ell}\coloneqq S\setminus S_{h}$. Fix an arbitrary ordering $\pi$ of $V(G)$. We start the algorithm with sets $T_{0}\coloneqq\varnothing$, $A_{0}\coloneqq V(G)\setminus S_{h}$, and the function $t_{0}(v)\coloneqq 0$ for every vertex $v\in V(G)$. For every integer $i\geqslant 0$, we pick a vertex $u_{i}\in A_{i}$ of the maximum degree in $G[A_{i}]$. In case there is more than one such vertex, we choose the one that comes earlier in $\pi$.
If $u_{i}\in S$, then let $T_{i+1}\coloneqq T_{i}\cup\{u_{i}\}$,
$t_{i+1}(v)\coloneqq\left\{\begin{array}[]{ll}t_{i}(v)+1&\textrm{if\ }v\in N_{G[A_{i}]}(u_{i}),\\ t_{i}(v)&\textrm{otherwise,}\end{array}\right.$
and
$A_{i+1}\coloneqq A_{i}\setminus(\{u_{i}\}\cup\{v\in A_{i}:t_{i+1}(v)>\Delta\}).$
If $u_{i}\notin S$, then let $T_{i+1}\coloneqq T_{i}$, $t_{i+1}(v)\coloneqq t_{i}(v)$ for every vertex $v\in V(G)$, and $A_{i+1}\coloneqq A_{i}\setminus\{u_{i}\}$. The algorithm terminates at step $K$, where $K$ equals the smallest $i$ such that $|A_{i}|\leqslant r$, and outputs
$S_{h},\quad T\coloneqq T_{K},\quad\text{and}\quad C\coloneqq A_{K}.$
Note that $|A_{i+1}|\leqslant|A_{i}|-1$ for every $i\geqslant 0$. Hence, the algorithm above indeed terminates after finitely many steps. Now, we verify that the outputs $S_{h}$, $T$, and $C$ satisfy all the claimed properties.
Claim 9.
For inputs $S^{\prime}$, $S^{\prime\prime}\subseteq V(G)$ where $|S^{\prime}|=|S^{\prime\prime}|$, if the algorithm outputs the same $S_{h}$ and $T$, then the algorithm outputs the same $C$.
Proof.
We write $\Delta^{\prime},S^{\prime}_{h},S^{\prime}_{\ell},T^{\prime}_{i},A^{\prime}_{i},t^{\prime}_{i},u^{\prime}_{i},K^{\prime},T^{\prime},C^{\prime}$ and $\Delta^{\prime\prime},S^{\prime\prime}_{h},S^{\prime\prime}_{\ell},T^{\prime\prime}_{i},A^{\prime\prime}_{i},t^{\prime\prime}_{i},u^{\prime\prime}_{i},K^{\prime\prime},T^{\prime\prime},C^{\prime\prime}$ for the corresponding parameters in the algorithm with input $S^{\prime}$ and $S^{\prime\prime}$, respectively. By assumption, we have $S^{\prime}_{h}=S^{\prime\prime}_{h}$ and hence, $A^{\prime}_{0}=A^{\prime\prime}_{0}$ and $u^{\prime}_{0}=u^{\prime\prime}_{0}$. Since $|S^{\prime}|=|S^{\prime\prime}|$, we also have $\Delta^{\prime}=\Delta^{\prime\prime}$. Besides, we have $T^{\prime}_{0}=T^{\prime\prime}_{0}=\varnothing$ and $t^{\prime}_{0}=t^{\prime\prime}_{0}$ by definition. Now, observe that if $C^{\prime}\neq C^{\prime\prime}$, then there is an index $i$ such that $u^{\prime}_{i}=u^{\prime\prime}_{i}$ but $u^{\prime}_{i}\in S^{\prime}$, $u^{\prime\prime}_{i}\notin S^{\prime\prime}$. Then, we have $u^{\prime}_{i}\in T^{\prime}$ but $u^{\prime\prime}_{i}\notin T^{\prime\prime}$, a contradiction to the assumption that $T^{\prime}=T^{\prime\prime}$. ∎
Claim 10.
$(S_{h}\cup T)\subseteq S$ and $S\setminus(S_{h}\cup T)\subseteq C$.
Proof.
By definition, we have $S_{h}\subseteq S$ and $T=T_{K}\subseteq S_{\ell}\subseteq S$, so $(S_{h}\cup T)\subseteq S$. For every vertex $v\in S\setminus(S_{h}\cup T)=S_{\ell}\setminus T_{K}$, if $v\notin C=A_{K}$, then there is an index $i<K$ such that $t_{i+1}(v)>\Delta$. This means that $v$ has more than $\Delta$ neighbors in $\{u_{1},\ldots,u_{i}\}\cap S$, so $v\in S_{h}$, a contradiction. ∎
Claim 11.
$|S_{h}|\leqslant 2\cdot\left(\frac{\mathrm{ex}(s,\,F)\log n}{d}\right)^{1/2}$.
Proof.
We have
$|S_{h}|\cdot\Delta\leqslant\sum_{v\in S}d_{S}(v)=2e(G[S])\leqslant 2\mathrm{ex}(s,F).$
Recalling that $\Delta=(\mathrm{ex}(s,F)\cdot d/\log n)^{1/2}$, we have
$|S_{h}|\leqslant\frac{2\cdot\mathrm{ex}(s,F)}{\Delta}=2\left(\frac{\mathrm{ex}(s,\,F)\log n}{d}\right)^{1/2}.\qed$
Claim 12.
$|T|\leqslant 32\cdot\left(\frac{\mathrm{ex}(s,\,F)\log n}{d}\right)^{1/2}$.
Proof.
For every positive integer $j\in[1,2\log n]$, let $I_{j}\coloneqq\{i\in[0,K-1]:n/2^{j}<|A_{i}|\leqslant n/2^{j-1}\}$. Let $J\coloneqq\{j:I_{j}\neq\varnothing\}$. Note that $\{I_{j}:j\in J\}$ forms a partition of $[0,K-1]$, as $2^{2\log n+1-1}>n$. For every $j\in J$, let $a_{j}$ and $b_{j}$ be the minimum and maximum elements in $I_{j}$, respectively. Since $T_{0}=\varnothing$, we have that $T_{K}$ can be expressed as the disjoint union
$T_{K}=\bigsqcup_{j\in J}\left(\left(T_{b_{j}}\setminus T_{a_{j}}\right)\cup\left(T_{b_{j}+1}\setminus T_{b_{j}}\right)\right),$
so
$|T|=|T_{K}|\leqslant 2\log n+\sum_{j\in J}|T_{b_{j}}\setminus T_{a_{j}}|.$ (2)
For every $j\in J$, we bound $|T_{b_{j}}\setminus T_{a_{j}}|$ by the observation
$\sum_{\begin{subarray}{c}i\,:\,a_{j}\leqslant i<b_{j}\\ u_{i}\in T_{K}\end{subarray}}d_{G[A_{i}]}(u_{i})\leqslant\sum_{v\in A_{a_{j}}}t_{b_{j}}(v),$ (3)
which holds because $t_{i+1}(v)$ increases by one for every $v\in N_{G[A_{i}]}(u_{i})\subseteq A_{a_{i}}$ when moving $u_{i}$ into $T_{i}$. For the left-hand side of (3), recall that $u_{i}$ is the vertex with maximum degree in $G[A_{i}]$, so
$d_{G[A_{i}]}(u_{i})\geqslant\frac{2e(G[A_{i}])}{|A_{i}|}\geqslant\frac{2e(G[A_{i}])}{|A_{a_{j}}|}.$
Note that for every $i<K$, we have $|A_{i}|>r$ by the terminating condition of our algorithm. By the assumption on $G$, we have $e(G[A_{i}])\geqslant d\cdot|A_{i}|^{2}\geqslant d\cdot|A_{b_{j}}|^{2}$. Hence,
$d_{G[A_{i}]}(u_{i})\geqslant\frac{2d|A_{b_{j}}|^{2}}{|A_{a_{j}}|}\geqslant\frac{2|A_{b_{j}}|}{|A_{a_{j}}|}\cdot d|A_{b_{j}}|\geqslant d|A_{b_{j}}|.$
For the right-hand side of (3), note that $t_{i}(v)$ is never larger than $\Delta+1$ for any $i$ and $v\in V(G)$. Therefore, by (3),
$|T_{b_{j}}\setminus T_{a_{j}}|\cdot d|A_{b_{j}}|\leqslant|A_{a_{j}}|\cdot(\Delta+1),$
which implies that
$|T_{b_{j}}\setminus T_{a_{j}}|\leqslant\frac{(\Delta+1)}{d}\cdot\frac{|A_{a_{j}}|}{|A_{b_{j}}|}\leqslant\frac{2(\Delta+1)}{d}\leqslant\frac{4\Delta}{d}.$
Note that $\log n\leqslant(\mathrm{ex}(s,\,F)\log n/d)^{1/2}$ by our assumption $\mathrm{ex}(s,F)\cdot d\geqslant\log n$. By (2), we have
$|T|=|T_{K}|\leqslant 2\log n+\frac{4\Delta}{d}\cdot 2\log n\leqslant 32\left(\frac{\mathrm{ex}(s,\,F)\log n}{d}\right)^{1/2}.\qed$
Claim 13.
$|C|\leqslant r$.
Proof.
The algorithm terminates only when $|C|=|A_{K}|\leqslant r$. ∎
Therefore, the sets $S_{h}$, $T$, and $C$ indeed satisfy all the claimed properties. ∎
Now, we prove Theorem 3 and use it to prove Theorems 1 and 4.
Proof of Theorem 3.
Let $c>0$ be sufficiently small and $C,N$ be sufficiently large. Without loss of generality, we may assume that $\mathrm{ex}(b,F)=\Omega(b)$. Recalling that $s=\left\lceil n^{\frac{\beta}{1-\alpha}}(\log n)^{\frac{3}{1-\alpha}}\right\rceil$ and $10\leqslant\alpha<$ 1, we have
$\mathrm{ex}\left(s,F\right)\cdot\delta n^{-\beta}=\Omega\left(s\cdot n^{-\beta}\right)=\Omega\left(n^{\frac{\beta}{1-\alpha}-\beta}(\log n)^{\frac{3}{1-\alpha}}\right)>\log n.$
Hence, we can apply Lemma 8 to the $F$-independent subsets $S\subseteq V(G)$ of size $s$ with parameters $d=\delta n^{-\beta}$ and $r=\gamma n^{\theta}$. Let
$t\coloneqq\left\lfloor 40\cdot\left({\mathrm{ex}(s,\,F)\log n}/{(\delta n^{-\beta})}\right)^{1/2}\right\rfloor,$
which is an upper bound for both $|S_{h}|$ and $|T|$ in Lemma 8, then
$t\log n=O\left(\left(n^{\frac{\beta}{1-\alpha}\cdot(1+\alpha)+\beta}(\log n)^{\frac{3}{1-\alpha}(1+\alpha)+1}\right)^{\frac{1}{2}}\right)\cdot\log n=O\left(n^{\frac{\beta}{1-\alpha}}(\log n)^{\frac{2+\alpha}{1-\alpha}}\right)\cdot\log n=O(s).$ (4)
Note that $s=o(n^{\theta})$ by the assumption that $\beta<\theta(1-\alpha)$. By Lemma 8, the number of $F$-independent subsets of $V(G)$ of size $s$ is at most
$\sum_{i=0}^{t}\sum_{j=0}^{t}\binom{n}{i}\binom{n}{j}\binom{\gamma n^{\theta}}{s}\leqslant(t+1)^{2}n^{2t}\binom{\gamma n^{\theta}}{s}\leqslant n^{4t}\left(\frac{e\gamma n^{\theta}}{s}\right)^{s}\leqslant e^{4t\log n}\left(\frac{e\gamma n^{\theta}}{s}\right)^{s}.$
Let $p\coloneqq 2c\cdot s/n^{\theta}$. We choose a random subset $B$ of $V(G)$, where every vertex is included with probability $p$, independently of each other. Let $\mathcal{F}$ be the (random) collection of $F$-independent subsets of $B$ of size $s$. Then, we have $\mathbb{E}[|B|]=np$ and
[…]
Let $B^{\prime}$ be the set obtained from $B$ by removing one vertex from every $S\in\mathcal{F}$ and $G^{\prime}\coloneqq G[B^{\prime}]$. Then, there is a choice of $G^{\prime}$ such that $G^{\prime}$ has no $F$-free subgraph of size $s$ and
$|V(G^{\prime})|\geqslant\mathbb{E}[|B^{\prime}|]\geqslant\mathbb{E}[|B|-|\mathcal{F}|]=\mathbb{E}[|B|]-\mathbb{E}[|\mathcal{F}|]\geqslant np-1=2csn^{1-\theta}-1\geqslant csn^{1-\theta}.$
Note that for $m\coloneqq csn^{1-\theta}$, we have $(1-\theta)\log n\leqslant\log m\leqslant\log n$ and
$m=csn^{1-\theta}\geqslant cn^{\frac{\beta}{1-\alpha}}(\log n)^{\frac{3}{1-\alpha}}\cdot n^{1-\theta}=cn^{\frac{(1-\alpha)(1-\theta)+\beta}{1-\alpha}}(\log n)^{\frac{3}{1-\alpha}},$
so
$s\leqslant 2n^{\frac{\beta}{1-\alpha}}(\log n)^{\frac{3}{1-\alpha}}\leqslant C\cdot m^{\frac{\beta}{(1-\alpha)(1-\theta)+\beta}}(\log m)^{\frac{3(1-\theta)}{(1-\alpha)(1-\theta)+\beta}}.\qed$
Proof of Theorem 1.
To prove the upper bound on $f_{F,K_{3}}(n)$, we use the following $K_{3}$-free graph given by Alon in [2].
Theorem 14 (Theorem 2.1 in [2]).
If $\ell$ is not divisible by $3$ and $m=2^{3\ell}$, then there exists a triangle-free $d_{\ell}$-regular graph $G_{\ell}$ on $m$ vertices, where $d_{\ell}=2^{\ell-1}(2^{\ell-1}-1)$, such that every eigenvalue $\nu$ of $G_{\ell}$, besides the largest, satisfies
$-9\cdot 2^{\ell}-3\cdot 2^{\ell/2}-1/4\leqslant\nu\leqslant 4\cdot 2^{\ell}+2\cdot 2^{\ell/2}+1/4.$
By the Expander Mixing Lemma (see for example Corollary 9.2.6 in [3]), we have that every set $X\subseteq V(G_{\ell})$ with $|X|\geqslant 10^{10}\cdot 2^{2\ell}$ spans at least $|X|^{2}/(10^{10}\cdot 2^{\ell})$ edges. Hence, we can apply Theorem 3 with $\beta=1/3$ and $\theta=2/3$ and get that there is a $K_{3}$-free graph $G^{\prime}_{\ell}$ with
$n=\Omega\left(m^{\frac{2-\alpha}{3-3\alpha}}(\log m)^{\frac{3}{1-\alpha}}\right)$
vertices and
$\alpha_{F}(G^{\prime}_{\ell})=O\left(m^{\frac{1}{3-3\alpha}}(\log m)^{\frac{3}{1-\alpha}}\right)=O\left(n^{\frac{1}{2-\alpha}}(\log n)^{\frac{3}{2-\alpha}}\right).$
Therefore, we have
$f_{F,K_{3}}(n)=O\left(n^{\frac{1}{2-\alpha}}(\log n)^{\frac{3}{2-\alpha}}\right).$
For $f_{F,K_{4}}(n)$, we use the following construction from [18].
Theorem 15 (Theorem 3 in [18]).
For every prime power $q\geqslant 2^{40}$, there exists a $K_{4}$-free graph $G_{q}$ with $m=q^{2}(q^{2}-q+1)$ vertices such that for every set $X$ of at least $2^{24}q^{2}$ vertices of $G_{q}$, we have $e(G_{q}[X])\geqslant{|X|^{2}}/({256q})$.
We apply Theorem 3 with $\beta=1/4$ and $\theta=1/2$ and get that there is a $K_{4}$-free graph $G_{q}^{\prime}$ with
$n=\Omega\left(m^{\frac{3-2\alpha}{4-4\alpha}}(\log m)^{\frac{3}{1-\alpha}}\right)$
vertices and
$\alpha_{F}(G_{q}^{\prime})=O\left(m^{\frac{1}{4-4\alpha}}(\log m)^{\frac{3}{1-\alpha}}\right)=O\left(n^{\frac{1}{3-2\alpha}}(\log n)^{\frac{6}{3-2\alpha}}\right).$
By Bertrand’s postulate that there is always a prime number between $a$ and $2a$ for every positive integer $a$, we conclude that
$f_{F,K_{4}}(n)=O\left(n^{\frac{1}{3-2\alpha}}(\log n)^{\frac{6}{3-2\alpha}}\right).\qed$
Proof of Theorem 4.
By the Expander Mixing Lemma, we have that there is $\gamma=\gamma(t)>0$ such that for every $A\subseteq V(G_{m})$ with $|A|\geqslant\gamma m^{1/2+1/(4t-6)}$, we have $e(A)\geqslant|A|^{2}\cdot\Omega(m^{-1/(2t-3)})$. Note that
$\frac{1}{2t-3}<\left(\frac{1}{2}+\frac{1}{4t-6}\right)\left(1-\alpha\right)$
by our assumption. The result follows by applying Theorem 3 with $\beta=1/(2t-3)$, and $\theta=1/2+1/(4t-6)$. ∎
3 Proof of Theorem 5
Following [12], we start from the following graph, which plays a crucial role in the proof of the lower bound on $r(4,t)$ [18] (see also [22]).
Lemma 16.
For every prime number $q$, there is a bipartite graph $F$ with vertex bipartition $(X,Y)$ such that
  1. (a)
    $|X|=q^{4}-q^{3}+q^{2}$, $|Y|=q^{3}+1$,
  2. (b)
    $d_{F}(x)=q+1$ for every $x\in X$ and $d_{F}(y)=q^{2}$ for every $y\in Y$,
  3. (c)
    $F$ is $C_{4}$-free,
  4. (d)
    there are no $\{x_{k}\}_{1\leqslant k\leqslant 4}\subseteq X$ and $\{y_{ij}\}_{1\leqslant i<j\leqslant 4}\subseteq Y$ such that $\{x_{i},y_{ij}\}$ and $\{x_{j},y_{ij}\}$ are edges in $F$ for every $1\leqslant i<j\leqslant 4$, i.e., $F$ does not contain the subdivision of $K_{4}$ as a subgraph with the four vertices of $K_{4}$ in $X$.
Let $r\geqslant 2$ be a fixed integer and $F$ be the graph from Lemma 16. For every vertex $y\in Y$, we randomly partition $N_{F}(y)$ into $r$ parts as $S_{1}(y)\cup\ldots\cup S_{r}(y)$. Let $H$ be the (random) graph on vertex set $X$, where $\{x_{1},x_{2}\}$ is an edge if and only if there exist a vertex $y\in Y$ and integers $1\leqslant i\neq j\leqslant r$ such that $x_{1}\in S_{i}(y)$ and $x_{2}\in S_{j}(y)$, i.e., for every $y\in Y$, we embed a complete $r$-partite graph in $N_{F}(y)$. Based on (c) and (d) in Lemma 16, the following is proved in [12]. See Lemma 2.2 and the discussion after it in [12].
Claim 17.
$H$ is always $K_{r+2}$-free.
As in the original proof of the lower bound on $r(4,t)$ in [18], the graph $H$ itself is not sufficient for proving Theorem 5, as $\alpha_{K_{s_{1},\ldots,s_{r}}}(H)$ may be too large. We need to do some further operations on $H$.
Recall that $s=\sum_{i=1}^{r}s_{i}$. We will consider the $s$-uniform hypergraph $\mathcal{H}$ on vertex set $X$, whose hyperedges consist of all the copies of $K_{s_{1},\ldots,s_{r}}$ in $H$. By definition, every $K_{s_{1},\ldots,s_{r}}$-independent set in $H$ is an independent set in $\mathcal{H}$, the number of which can be bounded by the following version of Hypergraph Container Lemma [5, 24]. See Corollary 2.8 in [12] for its proof.
Hereinafter, we denote $\{1,2,\ldots,m\}$ by $[m]$ for every positive integer $m$. For an $s$-uniform hypergraph $\mathcal{G}$ and $\ell\in[s]$, let $\Delta_{\ell}(\mathcal{G})$ be the maximum $\ell$-degree of $\mathcal{G}$, i.e., the maximum number of hyperedges containing an arbitrary fixed set of $\ell$ vertices.
Lemma 18.
For every positive integer $s$ and positive real numbers $p$ and $\lambda$, the following holds. Suppose that $\mathcal{G}$ is an $s$-uniform hypergraph such that $pv(\mathcal{G})$ and $v(\mathcal{G})/\lambda$ are integers, and for every $\ell\in[s]$,
$\Delta_{\ell}(\mathcal{G})\leqslant\lambda\cdot p^{\ell-1}\frac{e(\mathcal{G})}{v(\mathcal{G})}.$ (5)
Then, there exists a collection $\mathcal{C}$ of at most $v(\mathcal{G})^{spv(\mathcal{G})}$ sets of size at most $\left(1-\delta\lambda^{-1}\right)v(\mathcal{G})$ such that for every independent set $I$ in $\mathcal{G}$, there exists some $R\in\mathcal{C}$ with $I\subseteq R$, where $\delta=2^{-s(s+1)}$.
In order to guarantee that (5) in Lemma 18 is satisfied with a small value of $p$, we use the method from [12], choosing a set of copies of $K_{s_{1},\ldots,s_{r}}$ in $H$ by the following lemma.
Lemma 19 (Lemma 2.4 in [12]).
Let $q$ be a sufficiently large prime number. Then, with positive probability, for every $U\subseteq X$ such that $|U|\geqslant 500r^{2}q^{2}$, there exists some $\gamma\geqslant|U|/q^{2}$ such that the number of $y\in Y$ with $\gamma/(10r)\leqslant$ $\left|S_{i}(y)\cap U\right|\leqslant\gamma$ for every $i\in[r]$ is at least $|U|q/(8(\log q)\gamma)$.
We say that an instance of $H$ is good if it satisfies the conclusion of Lemma 19.
Lemma 20.
Let $q$ be a sufficiently large prime number and $a=q^{2-1/(s-1)}(\log q)^{3}$. If $H$ is good, then the number of sets $A\subseteq X$, where $|A|=a$ and $H[A]$ is $K_{s_{1},\ldots,s_{r}}$-free, is at most
$\left(q^{1/(s-1)}\right)^{a}.$
Proof.
Let $\mathcal{H}$ be the $s$-uniform hypergraph on vertex set $X$, whose hyperedges consist of all the copies of $K_{s_{1},\ldots,s_{r}}$ in $H$.
Claim 21.
For every $U\subseteq X$ with $|U|\geqslant 500s^{2}q^{2}$, there is a collection $\mathcal{C}_{U}\subseteq\mathcal{P}(U)$ of at most $q^{4sq^{2-1/(s-1)}}$ sets of size at most $(1-\Omega(\log q)^{-1})|U|$ such that for every independent set $I$ in $\mathcal{H}[U]$, there is some $R\in\mathcal{C}_{U}$ with $I\subseteq R$.
Proof.
By Lemma 19, there exist a real number $\gamma\geqslant|U|/q^{2}$ and a set $B\subseteq Y$ of vertices such that $|B|\geqslant|U|q/(8(\log q)\gamma)$ and $\gamma/(10r)\leqslant$ $\left|S_{i}(y)\cap U\right|\leqslant\gamma$ for every $y\in B$ and $i\in[r]$. Let $\mathcal{G}$ be the $s$-uniform hypergraph on vertex set $U$, where $\{x_{ij}:1\leqslant i\leqslant r,1\leqslant j\leqslant s_{i}\}$ is a hyperedge in $\mathcal{G}$ if and only if there is $y\in B$ such that $x_{ij}\in S_{i}(y)$ for every $i\in[r]$ and $j\in[s_{i}]$.
To apply Lemma 18 to $\mathcal{G}$, we need to choose proper values for $\lambda,p$ and verify that (5) is satisfied. Note that by (c) in Lemma 16, we have that $|N_{F}(y_{1})\cap N_{F}(y_{2})|\leqslant 1$ for distinct $y_{1},y_{2}\in B$. Hence, by the definition of $\mathcal{G}$, we have
$e(\mathcal{G})\geqslant|B|\cdot\prod_{i=1}^{r}\binom{|S_{i}(y)\cap U|}{s_{i}}=\frac{|U|q}{8(\log q)\gamma}\cdot\Omega\left(\left(\frac{\gamma}{10r}\right)^{s}\right)=\Omega\left(\frac{|U|q\gamma^{s-1}}{\log q}\right),$
which implies that
$\frac{e(\mathcal{G})}{v(\mathcal{G})}=\Omega\left(\frac{q\gamma^{s-1}}{\log q}\right).$
For every $x\in U$, there can be at most $d_{F}(x)=q+1$ vertices $y\in B$ such that $x\in N_{F}(y)$, so
$\Delta_{1}(\mathcal{G})=(q+1)\cdot O\left(\gamma^{s-1}\right).$
For every set of $\ell$ $(2\leqslant\ell\leqslant s)$ vertices $\{x_{1},\ldots,x_{\ell}\}\subseteq U$, there can be at most one vertex $y\in B$ such that $\{x_{1},\ldots,x_{\ell}\}\subseteq N_{F}(y)$, so
$\Delta_{\ell}(G)=O\left(\gamma^{s-\ell}\right).$
Let $C$ be a sufficiently large constant compared to $s$ and $r$. Let $\lambda$ be a real number between $C\log q$ and $2C\log q$ such that $v(\mathcal{G})/\lambda$ is an integer. Let $p$ be a real number between $(\gamma q^{1/(s-1)})^{-1}$ and $2(\gamma q^{1/(s-1)})^{-1}$ such that $pv(\mathcal{G})$ is an integer. Then, (5) is satisfied. By Lemma 18, there exists a collection $\mathcal{C}_{U}$ of at most
$v(G)^{spv(G)}\leqslant|U|^{s\cdot 2(\gamma q^{1/(s-1)})^{-1}\cdot|U|}\leqslant(q^{4})^{s\cdot 2\left(\frac{|U|}{q^{2}}\cdot q^{1/(s-1)}\right)^{-1}\cdot|U|}\leqslant q^{8sq^{2-1/(s-1)}}$
sets of size at most
$\left(1-2^{-s(s+1)}\lambda^{-1}\right)v(\mathcal{G})=\left(1-\Omega\left((\log q)^{-1}\right)\right)|U|$
such that for every independent set $I$ in $\mathcal{G}$, there exists some $R\in\mathcal{C}_{U}$ with $I\subseteq R$. The conclusion in 21 follows by noting that $\mathcal{G}$ is a subhypergraph of $\mathcal{H}$, so every independent set in $\mathcal{H}[U]$ is an independent set in $\mathcal{G}$. ∎
Claim 22.
There is a collection $\mathcal{C}\subseteq\mathcal{P}(X)$ of at most $q^{O(q^{2-1/(s-1)}(\log q)^{2})}$ sets of size at most $500s^{2}q^{2}$ such that for every independent set $I$ in $\mathcal{H}$, there is some $R\in\mathcal{C}$ such that $I\subseteq R$.
Proof.
We prove by induction that for every positive integer $j$, there is a collection $\mathcal{C}^{(j)}$ of at most $q^{8jsq^{2-1/(s-1)}}$ sets of size at most $\max\left\{500s^{2}q^{2},\,(1-\Omega((\log q)^{-1}))^{j}|X|\right\}$ such that for every independent set $I$ in $\mathcal{H}$, there is some $R\in\mathcal{C}^{(j)}$ such that $I\subseteq R$. 22 then follows by letting $j$ be a suitable integer of order $\Theta((\log q)^{2})$.
The case $j=1$ holds by applying 21 with $U=X$. Now, suppose we have a proper collection $\mathcal{C}^{(j)}$. For every $U\in\mathcal{C}^{(j)}$ with $|U|>500s^{2}q^{2}$, we apply 21 to $U$ and get a collection $\mathcal{C}_{U}$. Let
$\mathcal{C}^{(j+1)}\coloneqq\left\{U\in\mathcal{C}^{(j)}:|U|\leqslant 500s^{2}q^{2}\right\}\cup\bigcup_{U\in\mathcal{C}^{(j)}:|U|>500s^{2}q^{2}}\mathcal{C}_{U}.$
By the induction hypothesis, we have
$\mathcal{C}^{(j+1)}\leqslant|\mathcal{C}^{(j)}|\cdot q^{8sq^{2-1/(s-1)}}\leqslant q^{8jsq^{2-1/(s-1)}}\cdot q^{8sq^{2-1/(s-1)}}=q^{8(j+1)sq^{2-1/(s-1)}}.$
For every set $R\in\mathcal{C}^{(j+1)}$, either $|R|\leqslant 500s^{2}q^{2}$ or there is $U\in\mathcal{C}^{(j)}$ such that $R\in\mathcal{C}_{U}$, in which case we have
$|R|\leqslant(1-\Omega((\log q)^{-1}))|U|\leqslant(1-\Omega((\log q)^{-1}))^{j+1}|X|.$
So, $|R|\leqslant\max\left\{500s^{2}q^{2},\,(1-\Omega((\log q)^{-1}))^{j+1}|X|\right\}$. For every independent set $I\in\mathcal{H}$, there is some $U\in\mathcal{C}^{(j)}$ such that $I\subseteq U$. If $|U|\leqslant 500s^{2}q^{2}$, then $I\subseteq U\in\mathcal{C}^{(j+1)}$; if $|U|>500s^{2}q^{2}$, then there is some $R\in\mathcal{C}_{U}$ such that $I\subseteq R\in\mathcal{C}_{U}\subseteq\mathcal{C}^{(j+1)}$. Therefore, $C^{(j+1)}$ indeed has the properties claimed. ∎
By 22, the number of independent sets of size $a=q^{2-1/(s-1)}(\log q)^{3}$ in $\mathcal{H}$ is at most
$q^{O(q^{2-1/(s-1)}(\log q)^{2})}\cdot\binom{500s^{2}q^{2}}{a}\leqslant q^{O(q^{2-1/(s-1)}(\log q)^{2})}\cdot\left(\frac{500e\cdot s^{2}q^{2}}{a}\right)^{a}\leqslant\left(q^{1/(s-1)}\right)^{a}.\qed$
Proof of Theorem 5.
Let $q$ be a sufficiently large prime number and fix a good $H$, whose existence is guaranteed by Lemma 19. Note that $H$ is $K_{r+2}$-free by 17. Let $a\coloneqq q^{2-1/(s-1)}(\log q)^{3}$ and $\mathcal{A}$ be the collection of sets $A\subseteq X$, where $|A|=a$ and $H[A]$ is $K_{s_{1},\ldots,s_{r}}$-free. By Lemma 20, we have $|\mathcal{A}|\leqslant\left(q^{1/(s-1)}\right)^{a}$.
We choose a random subset $X^{\prime}$ of $X$, where every vertex in $X$ is kept independently with probability $q^{-1/(s-1)}$. Let $H^{\prime}\coloneqq H[X^{\prime}]$ and $\mathcal{A}^{\prime}$ be the (random) collection of sets $\{A\in\mathcal{A}:A\subseteq X^{\prime}\}$. Let $H^{\prime\prime}$ be the (random) graph obtained from $H^{\prime}$ by removing one vertex in every $A\in\mathcal{A}^{\prime}$. Note that $H^{\prime\prime}$ is also $K_{r+2}$-free and does not contain any $K_{s_{1},\ldots,s_{r}}$-free set of size $a$. The expected number of vertices in $H^{\prime\prime}$ is at least
$\mathbb{E}[|X^{\prime}|-|\mathcal{A}^{\prime}|]=\mathbb{E}[|X^{\prime}|]-\mathbb{E}[|\mathcal{A}^{\prime}|]\geqslant|X|\cdot q^{-1/(s-1)}-1\geqslant\frac{1}{2}q^{4-1/(s-1)}.$
Therefore, for every sufficiently large prime number $q$, there exists a $K_{r+2}$-free graph $H^{\prime\prime}$ with at least $\frac{1}{2}q^{4-1/(s-1)}$ vertices and $\alpha_{K_{s_{1},\ldots,s_{r}}}(G)<a=q^{2-1/(s-1)}(\log q)^{3}$. By Bertrand’s postulate that there is always a prime number between $a$ and $2a$ for every positive integer $a$, we conclude that
$f_{K_{s_{1},\ldots,s_{r}},K_{r+2}}(n)=O\left(n^{\frac{2s-3}{4s-5}}(\log n)^{3}\right).\qed$
4 Remark on $\mathrm{ex}(G(n,p),C_{4})$
The previous bounds on $\mathrm{ex}(G(n,p),C_{4})$ are the following.
Theorem 23 ([11, 14, 19]).
For every $\delta>0$, we have
[…]
Proposition 7 improves the bounds in the first two ranges as follows:
[…]
Note that the “construction” for the lower bound is the same, and we extend the range of $p$, where it is best possible. In order to obtain the claimed improvement, we use the following theorem and the first moment method. Denote $f(C_{4},n,T)$ the number of (labelled) $C_{4}$-free graphs on $n$ vertices with $T$ edges.
Theorem 24 (Lemma 5 in [14]).
Let $c=(2^{65}/3)^{1/3}$ and $\varepsilon=3/(13\cdot 2^{65})$. For every integer $T>cn^{4/3}(\log n)^{1/3}$, we have
$f(C_{4},n,T)<\left(\frac{4en^{3}}{T^{2}}\right)^{T}\exp\left(\varepsilon^{-1}\left(n\log n\right)^{4/3}\right).$
Proof of Proposition 7.
Let $C$ be a sufficiently large real number and $T$ be an integer between $Cp^{1/2}n^{3/2}$ and $2Cp^{1/2}n^{3/2}$. Let $c=(2^{65}/3)^{1/3}$ and $\varepsilon=3/(13\cdot 2^{65})$, as in Theorem 24. Note that
$T\geqslant Cp^{1/2}n^{3/2}\geqslant Cn^{-1/6}(\log n)^{4/3}\cdot n^{3/2}=Cn^{4/3}(\log n)^{4/3}>cn^{4/3}(\log n)^{1/3}.$ (6)
So, by Theorem 24, the expected number of $C_{4}$-free subgraphs with $T$ edges in $G(n,p)$ is at most
$p^{T}\left(\frac{4en^{3}}{T^{2}}\right)^{T}\exp\left(\varepsilon^{-1}\left(n\log n\right)^{4/3}\right)$ $\leqslant\left(\frac{4epn^{3}}{C^{2}pn^{3}}\right)^{T}\exp\left(\varepsilon^{-1}\left(n\log n\right)^{4/3}\right)$
$=\left(\frac{4e}{C^{2}}\right)^{T}\exp\left(\varepsilon^{-1}\left(n\log n\right)^{4/3}\right)=o(1).$
Therefore, with high probability, there is no $C_{4}$-free subgraph with $T$ edges in $G(n,p)$ and hence, we have $\mathrm{ex}(G(n,p),C_{4})<T$. ∎
Acknowledgment
We thank the anonymous referees for many helpful comments that improved the quality of the manuscript. The authors are grateful to Noga Alon, Sam Mattheus, Dhruv Mubayi, Jacques Verstraëte, and Ethan White for helpful discussions.